A 40 kg box is being moved by a man and a woman over a rough horizontal surface (u = 0.4). The man pulls up
on a box with a rope with a force of 500 N [40 º above the horizontal] and the woman pushes the box with a
force of 150 N directly horizontally (same direction as the man). The box moves 30 m.
a. Calculate the work done by each force acting on the box.
b. What is the change in energy for the object?
1
Expert's answer
2020-05-15T08:42:59-0400

a)


"W_M=F_Md\cos{40}=(500)(30)\cos{40}=11500\ J"

"W_W=F_Wd=(150)(30)=4500\ J"

"W_f=-\mu mgd=-0.4(40)(9.8)(30)=-4700\ J"

b)


"\Delta E=W_W+W_M+W_f\\=11500+4500-4700=11300\ J"